OpenStudy (anonymous):

MEDAL REWARD!! :D Working together, two pumps can drain a certain pool in 5 hours. If it takes the older pump 15 hours to drain the pool by itself, how long will it take the newer pump to drain the pool on its own? Do not do any rounding.

OpenStudy (anonymous):

Let t = time required by new pump to drain the pool Let the completed job (drained pool) = 1 \[\frac{ 5 }{ 15 }+\frac{ 5 }{ t }=1\] Multiply by 15t t(5) + 15(5) = 15 5t + 75 = 15t 75 = 15t - 5t 75 = 10t t = 75/10 t = 7.5 hrs new pump alone Check this 5/15 + 5/7.5 = .33 + .67 = 1

OpenStudy (uri):